Q)A train stopping at two staions 4 km apart takes 4 min on the journey from one of the station to other.Assuming that it first accelerates with a uniform acceleration x and then that of uniform retardation y, prove that 1/x + 1/y = 2.
PLZ SOLVE IT SIR............

Grade:11

1 Answers

Abhishek Kumar
9 years ago
Let it accelerates for time t,
v = xt
Then it decelerates for (240-t)
0 – v = y(240-t)
If you draw the v-t graph, it is a triangle with one side made by a straight line from origin to v, with slope x. Then v goes down to zero at t= 240s. Total distance travelled = Area under v-t = ½.240.v = 4000